OpenStudy (anonymous):

David has 4 1/3 feet of board for a workshop project. He needs 2/3 of a foot for each piece of the project. How many complete pieces of 2/3 foot board will he be able to cut? Answer 8 2 3 6

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

@privetek

OpenStudy (anonymous):

@EvrydayAzn

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

@JakeV8 I got 8

OpenStudy (anonymous):

Ok, I'm just now reading it... how did you get 8?

OpenStudy (anonymous):

I mutiply cross right?

OpenStudy (anonymous):

First, I would make 4 1/3 into an improper fraction... 4 is 12/3, right? so 4 1/3 is = 13/3 And you want to know how many FULL pieces of 2/3 length come out of that 13/3... So you want to divide: 14/3 divided by 2/3\[\frac{ 14 }{ 3 } \div \frac{ 2 }{ 3 } = \frac{ 14 }{ 3 } \times \frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

sorry, I messed up that equation... wrote "14" instead of "13" in both places :(

OpenStudy (anonymous):

This is correct:\[\frac{ 13 }{ 3 } \div \frac{ 2 }{ 3 } = \frac{ 13 }{ 3 } \times \frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

39/8

OpenStudy (anonymous):

the bottom is 6, not 8. But rather than do that, cancel the 3 on top with the 3 on bottom... you get 13/2

OpenStudy (anonymous):

So B is the correct answer.

OpenStudy (anonymous):

No? If you can cut 13/2 pieces of wood, that means you are cutting 6.5 pieces. But that last half piece doesn't count... you are asked how many FULL pieces you can cut. So it's only 6 full pieces.